### v1

@@ 2 Related work
Object-centric slot representations have been studied extensively in the literature [22, 28, 8, 77, 21, 42, 27, 26, 25, 7]. Often, these representations are expected to represent predefined objects and attributes, and evaluated in small-scale or synthetic settings. Recently, the use of slot attention on features produced by a learned encoder [60, 53, 1], and utilizing unsupervised saliency masks for object discovery [75] has enabled unsupervised learning of object-centric representations in larger scale settings. Related to these ideas, Sparo attends to features produced by a transformer using a single attention head per slot, and each head is capable of learning saliency maps without supervision [9]. However, we do not constrain our slots to correspond to predefined types of objects in the data, nor do we impose any independence conditions between the slots. Self-supervised simplicial embeddings (SEM) [38], with only a soft-discrete slot structure constraint, enable improved generalization on downstream tasks. Sparo replaces the separation of $\mathrm{softmax}$ application between SEM slots with a separation of underlying attention operations.
@@ 3.2 Separate-head attention read-out (Sparo)
We replace the pooling operation of transformer encoders with a concatenation of outputs of $L$ single-head attention mechanisms. Our module, Sparo, acts on the transformer outputs ${\bm{H}}=\{{\bm{h}}_{1},\ldots,{\bm{h}}_{n}\}\in\mathbb{R}^{n\times d}$ to produce the encoder output ${\bm{y}}=\mathrm{concat}({\bm{y}}_{1},\ldots,{\bm{y}}_{L})\in\mathbb{R}^{LV}$, where Sparo slot ${\bm{y}}_{l}\in\mathbb{R}^{V}$, $l\in[L]$. We illustrate the architecture of Sparo in Fig. 1. Each Sparo slot is produced as follows:
@@ 3.3 CLIP with Sparo
We add Sparo to both the image and text encoders of CLIP [54], setting the same values of $L$ and $V$ in both encoders. Each Sparo slot is $\ell_{2}$-normalized separately, and we divide their concatenation by $\sqrt{L}$ to produce the global $\ell_{2}$-normalized encoding. With this scheme, the cosine similarity $\mathrm{sim}({\bm{y}}^{\text{i}},{\bm{y}}^{\text{t}})$ between the image encoding ${\bm{y}}^{\text{i}}$ and the text encoding ${\bm{y}}^{\text{t}}$ becomes the expected cosine similarity between corresponding slot encodings:
$\mathrm{sim}({\bm{y}}^{\text{i}},{\bm{y}}^{\text{t}})={{\bm{y}}^{\text{i}}}^{\top}{\bm{y}}^{\text{t}}=\sum_{l=1}^{L}\lVert{\bm{y}}^{\text{i}}_{l}\rVert\lVert{\bm{y}}^{\text{t}}_{l}\rVert\frac{{{\bm{y}}^{\text{i}}_{l}}^{\top}{\bm{y}}^{\text{t}}_{l}}{\lVert{\bm{y}}^{\text{i}}_{l}\rVert\lVert{\bm{y}}^{\text{t}}_{l}\rVert}=\sum_{l=1}^{L}\frac{1}{L}\mathrm{sim}({\bm{y}}^{\text{i}}_{l},{\bm{y}}^{\text{t}}_{l})=\operatorname*{\mathbb{E}}_{l\sim[L]}\left[\mathrm{sim}({\bm{y}}^{\text{i}}_{l},{\bm{y}}^{\text{t}}_{l})\right].$ (2)
We use the standard CLIP loss to train the model, which maximizes $\mathrm{sim}({\bm{y}}^{\text{i}},{\bm{y}}^{\text{t}})$ of encodings ${\bm{y}}^{\text{i}}$ and ${\bm{y}}^{\text{t}}$ for aligned image-text pairs and minimizes it across all other pairings. With Sparo’s separate-head attention bottleneck, maximizing $\operatorname*{\mathbb{E}}_{l\sim[L]}\left[\mathrm{sim}({\bm{y}}^{\text{i}}_{l},{\bm{y}}^{\text{t}}_{l})\right]$ for aligned pairs across the training dataset encourages both modalities to learn similar attention semantics per slot $l\in[L]$. Consequently, Sparo imparts a prior that both modalities encode a shared world with $L$ selectively attendable concepts, and similarities between inputs can be expressed as their average similarity for these concepts.
@@ 4 Results
Over Secs. 4 and 5, we evaluate our method on a variety of downstream tasks, analyze Sparo’s representation structure with the ability to intervene on selected slots, perform ablations to support our design, illustrate the effect of changing Sparo’s hyperparameters, and visualize example learned concepts.
@@
For all experiments in this section, we train CLIP models using the open-source OpenCLIP [34] project. We consider two model sizes: those with a ViT-B/16 visual backbone (CLIP16) and those with a ViT-B/32 backbone (CLIP32). For models trained with Sparo, we replace the last transformer block of the base transformer backbones. For our chosen settings, this ensures that the resulting model size is comparable to the original model size. Further, since only the CLS token output from the image transformer and the EOS token output from the text transformer are used for the standard CLIP encodings, removing one transformer block ensures that we do not gain an unfair advantage by attending to positions from a layer that standard CLIP discards. Note that standard CLIP also retains the MLP as a part of its final transformer block, which is absent in the Sparo module. We also compare our encodings with those produced by CLIP with global average pooling (CLIP+GAP), which enjoys the advantage of being able to use all of the final layer outputs. When using Sparo, we use $L=V=64$ when training on Conceptual Captions, and $L=128,V=64$ when training on LAION-400M.
@@ 4.1 Zero-shot recognition, robustness, and compositionality
We evaluate the zero-shot classification accuracy of trained CLIP models on ImageNet [17] and a set of robustness benchmarks including ImageNet-V2 [57], ImageNet-Sketch [68], ImageNet-A [31], ImageNet-R [30], and ObjectNet [4]. Additionally, to evaluate the compositionality of our learned encodings in terms of objects, attributes, and relations, we test our trained models on SugarCrepe [33]. We report our results in Tab. 1, and provide more fine-grained SugarCrepe numbers in Sec. 0.B.1. We find that using Sparo encodings outperforms CLIP and CLIP+GAP in all settings considered. The improved results on the SugarCrepe benchmark suggests that Sparo encodings exhibit compositionality at a level that enable better binding of object, attribute, and relation properties than baseline encodings on average.
@@
Finally, we also use the VTAB benchmark [76] to evaluate Sparo on a diverse set of datasets [23, 36, 15, 49, 51, 69, 74, 29, 14, 67, 20, 35, 47, 39, 5, 24] that vary more significantly from the training data. We present the relative zero-shot classification accuracy improvements of CLIP+Sparo over CLIP+GAP in Fig. 2. We see that Sparo outperforms standard CLIP encodings for a majority of VTAB tasks in all four settings.
@@ 5 Analysis
In this section, we provide insights into encodings learned by Sparo. We demonstrate the ability of using Sparo’s slot structure to perform post-hoc selection of slots for improving downstream performance. We then show that without this structure, such selection is more prone to overfitting. Next, we validate the design choice of using single-head attention to model the notion of selective attention through controlled ablations. We analyze the effect of Sparo’s hyperparameters on performance and model size. Finally, we visualize example concepts that CLIP models trained with Sparo learn to attend to.
@@ 5.2 Robustness of Sparo’s slot structure
5.3 Ablating the bottleneck and cross-modal alignment of Sparo heads
@@
Model Replace last block Cross- attention Slot-wise LayerNorm Slot-wise projection Params ImageNet zero-shot
CLIP32 (C) N/A N/A N/A N/A 151M 0.329
$\mathcal{C}$+GAP 151M 0.345
$\mathcal{C}$+AttPool [72] No Multi-head No No 440M 0.312
Ablated $\mathcal{C}$+Sparo Yes 306M 0.344
Yes No 440M 0.338
Yes 306M 0.359
Separate-head No No 306M 0.344
Yes 172M 0.358
Yes No 306M 0.344
Yes 172M 0.374
$\mathcal{C}$+Sparo N/A N/A 161M 0.372
Yes 151M 0.370
@@
We present our results in Tab. 6, showing a clear advantage of Sparo over AttnPool. Additionally, both the choices of separate-head attention over multi-head and slot-wise operations over the alternatives result in improved generalization and a reduced model size. While the former supports the effectiveness of the separate-head attention bottleneck, the latter validates the benefits of the prior for shared separately-attendable concepts between the modalities enabled by Sparo.
5.4 Hyperparameters $L$ and $V$
(a) ImageNet zero-shot accuracy.
(b) Number of model parameters.
Figure 3: Effect of varying $L$ and $V$ values for CLIP32+Sparo, and the embedding size for CLIP32 and CLIP32+GAP, when training on CC15M.
Sparo encodings are concatenations of $L$ separately-attended slots of $V$ dimensions each. We study the effect of $L$ and $V$ with CLIP32+Sparo trained on CC15M, comparing its ImageNet zero-shot accuracy and model size with that of CLIP32 and CLIP32+GAP in Fig. 3. Generally, performance is poor for very low values of $L$ and $V$, and improves when they are increased. Improvements remain monotonic with $L$ in our experimental range, but the returns diminish for larger values of $L$. For each increment of $L$, we add additional ${\bm{q}}_{l}$ and ${\bm{K}}_{l}$ parameters to the model, and considering the trade-off between model size and performance gains becomes vital. Unlike $L$, increasing $V$ beyond optimal values leads to a performance drop. With very large values of $V$, we produce low-rank slots that can potentially encode arbitrary harmful biases for downstream tasks.
@@ 5.5 Visualizations
To ease our interpretation of attended concepts, we choose concepts to visualize that have sharp attention in the text modality, attend to non-overlapping text tokens, and have cosine similarity more than 0.75 with the corresponding image slots. We present examples of these filtered concepts in Fig. 4. In this example, we see that Sparo is capable of attending to different concepts such as subject of the scene, the activity represented, and the surrounding location, for both vision and text. We provide additional visualizations in Sec. 0.B.3.
@@ 6 Discussion
Humans utilize a prior for compositionality to coherently represent different subsets of salient aspects of the world. We can enrich our comprehension by selectively attending to new aspects or specialize to a task by filtering out distracting ones. We introduce Sparo with a goal of imparting a similar prior to transformers in representation learning frameworks by partitioning encodings into separately-attended concepts. Although we see evidence of disentanglement through positive results of post-hoc concept selection and visualization of attention maps, we do not impose any explicit independence or disentanglement constraints. In our representation learning frameworks, the pressure for learning distinct concepts arises from the need to explain the factors of variation through the training objective while remaining bottlenecked to represent features that can be accumulated by a single head of attention with embedded queries and limited dimensionality. However, auxiliary training objectives to impose stronger distributional conditions on the learned concepts can be explored as promising future directions. Beyond pretraining, we highlight that more sophisticated post-hoc concept selection approaches than explored in our work, like using human interaction or set cover algorithms, can further improve the downstream utility of Sparo.
@@ Appendix 0.B Additional results
0.B.1 Fine-grained SugarCrepe performance
Table 0.B.1: Fine-grained SugarCrepe zero-shot accuracies for CLIP models trained on Conceptual Captions and LAION-400M.

### v2

@@ 2 Related work
Object-centric slot representations have been studied extensively in the literature [22, 28, 8, 79, 21, 43, 27, 26, 25, 46, 7]. Often, these representations are expected to represent predefined objects and attributes, and evaluated in small-scale or synthetic settings. Recently, the use of slot attention on features produced by a learned encoder [62, 55, 1], and utilizing unsupervised saliency masks for object discovery [77] has enabled unsupervised learning of object-centric representations in larger scale settings. Related to these ideas, Sparo attends to features produced by a transformer using a single attention head per slot, and each head is capable of learning saliency maps without supervision [9]. However, we do not constrain our slots to correspond to predefined types of objects in the data, nor do we impose any independence conditions between the slots. Self-supervised simplicial embeddings (SEM) [39], with only a soft-discrete slot structure constraint, enable improved generalization on downstream tasks. Sparo replaces the separation of $\mathrm{softmax}$ application between SEM slots with a separation of underlying attention operations.
@@ 3.2 Separate-head attention read-out (Sparo)
We replace the pooling operation of transformer encoders with a concatenation of outputs of $L$ single-head attention mechanisms. Sparo acts on the transformer outputs ${\bm{H}}=\{{\bm{h}}_{1},\ldots,{\bm{h}}_{n}\}\in\mathbb{R}^{n\times d}$ to produce the encoder output ${\bm{y}}=\mathrm{concat}({\bm{y}}_{1},\ldots,{\bm{y}}_{L})\in\mathbb{R}^{LV}$, where Sparo slot ${\bm{y}}_{l}\in\mathbb{R}^{V}$, $l\in[L]$. We illustrate the architecture of Sparo in Fig. 1. Each Sparo slot is produced as:
@@ 3.3 CLIP with Sparo
We add Sparo to both the image and text encoders of CLIP [56], setting the same values of $L$ and $V$ in both encoders. Each Sparo slot is $\ell_{2}$-normalized separately, and we divide their concatenation by $\sqrt{L}$ to produce the global $\ell_{2}$-normalized encoding. With this scheme, the cosine similarity between the image encoding ${\bm{y}}^{\text{i}}$ and the text encoding ${\bm{y}}^{\text{t}}$ becomes the expected cosine similarity between corresponding slot encodings:
${{\bm{y}}^{\text{i}}}^{\top}{\bm{y}}^{\text{t}}=\sum_{l=1}^{L}\lVert{\bm{y}}^{\text{i}}_{l}\rVert\lVert{\bm{y}}^{\text{t}}_{l}\rVert\frac{{{\bm{y}}^{\text{i}}_{l}}^{\top}{\bm{y}}^{\text{t}}_{l}}{\lVert{\bm{y}}^{\text{i}}_{l}\rVert\lVert{\bm{y}}^{\text{t}}_{l}\rVert}=\sum_{l=1}^{L}\frac{1}{L}\mathrm{sim}({\bm{y}}^{\text{i}}_{l},{\bm{y}}^{\text{t}}_{l})=\operatorname*{\mathbb{E}}_{l\sim[L]}\left[\mathrm{sim}({\bm{y}}^{\text{i}}_{l},{\bm{y}}^{\text{t}}_{l})\right].$ (2)
We train the model using the standard CLIP loss which maximizes $\mathrm{sim}({\bm{y}}^{\text{i}},{\bm{y}}^{\text{t}})$ of encodings ${\bm{y}}^{\text{i}}$ and ${\bm{y}}^{\text{t}}$ for aligned image-text pairs and minimizes it across all other pairings. With Sparo’s separate-head attention bottleneck, maximizing $\operatorname*{\mathbb{E}}_{l\sim[L]}\left[\mathrm{sim}({\bm{y}}^{\text{i}}_{l},{\bm{y}}^{\text{t}}_{l})\right]$ for aligned pairs across the training dataset encourages both modalities to learn similar attention semantics per slot $l\in[L]$. Consequently, Sparo imparts a prior that both modalities encode a shared world with $L$ selectively attendable concepts, and similarities between inputs can be expressed as their average similarity for these concepts.
@@ 4 Results
In Secs. 4 and 5, we evaluate our method on a variety of downstream tasks, analyze Sparo’s representation structure with the ability to intervene on selected slots, perform ablations to support our design, and visualize the learned concepts.
@@
For all experiments in this section, we train CLIP models using the open-source OpenCLIP [35] project. We consider two model sizes: those with a ViT-B/16 visual backbone (CLIP16) and those with a ViT-B/32 backbone (CLIP32). For models trained with Sparo, we replace the last transformer block of the base transformer backbones. For our chosen settings, this ensures that the resulting model size is comparable to the original model size. Further, since only the CLS token output from the image transformer and the EOS token output from the text transformer are used for the standard CLIP encodings, removing one transformer block ensures that we do not gain an unfair advantage by attending to positions from a layer that standard CLIP discards. Note that standard CLIP also retains the MLP as a part of its final transformer block, which is absent in the Sparo module. We also compare our encodings with those produced by CLIP with global average pooling (CLIP+GAP), which enjoys the advantage of being able to use all of the final layer outputs. When using Sparo, we use $L=V=64$ when training on Conceptual Captions, and $L=128,V=64$ when training on LAION-400M. We also provide results for CLIP with residual network (ResNet) [29] encoders in Sec. 0.B.3.
@@ 4.1 Zero-shot recognition, robustness, and compositionality
We evaluate the zero-shot classification accuracy of trained CLIP models on ImageNet [17] and a set of robustness benchmarks including ImageNet-V2 [59], ImageNet-Sketch [70], ImageNet-A [32], ImageNet-R [31], and ObjectNet [4]. Additionally, to evaluate the compositionality of our learned encodings in terms of objects, attributes, and relations, we test our trained models on SugarCrepe [34]. We report our results in Tab. 1, and provide more fine-grained SugarCrepe numbers in Sec. 0.B.4. We find that using Sparo encodings outperforms CLIP and CLIP+GAP in all settings considered. The improved results on the SugarCrepe benchmark suggests that Sparo encodings exhibit compositionality at a level that enable better binding of object, attribute, and relation properties than baseline encodings on average.
@@
Finally, we also use the VTAB benchmark [78] to evaluate Sparo on a diverse set of datasets [23, 37, 15, 51, 53, 71, 76, 30, 14, 69, 20, 36, 49, 40, 5, 24] that vary more significantly from the training data. We present the relative zero-shot classification accuracy improvements of CLIP+Sparo over CLIP+GAP in Fig. 2 and the corresponding absolute values in Sec. 0.B.6. We see that Sparo outperforms standard CLIP encodings for a majority of VTAB tasks in all four settings.
@@ 5 Analysis
In this section, we provide insights into encodings learned by Sparo. We demonstrate the ability of using Sparo’s slot structure to perform post-hoc selection of slots for improving downstream performance. We then show that without this structure, such selection is more prone to overfitting. Next, we validate the design choice of using single-head attention to model the notion of selective attention through controlled ablations. Finally, we visualize example concepts that CLIP models trained with Sparo learn to attend to.
@@ 5.2 Robustness of Sparo’s slot structure
5.3 Ablating the bottleneck and cross-modal alignment of Sparo
@@
Model Replace last block Cross- attention Slot-wise LayerNorm Slot-wise projection Params flops (×$\times 10^{9}$) ImageNet zero-shot
CLIP32 (C) N/A N/A N/A N/A 151M 7.4 0.329
$\mathcal{C}$+GAP 151M 7.4 0.345
$\mathcal{C}$+AttPool No Multi-head No No 440M 8.9 0.312
Ablated $\mathcal{C}$+Sparo Yes 306M 8.8 0.344
Yes No 440M 8.9 0.338
Yes 306M 8.8 0.359
Separate-head No No 306M 8.8 0.344
Yes 172M 8.7 0.358
Yes No 306M 8.8 0.344
Yes 172M 8.7 0.374
$\mathcal{C}$+Sparo N/A N/A 161M 8.0 0.372
Yes 151M 7.4 0.370
@@
We present our results in Tab. 6, showing a clear advantage of Sparo over AttnPool. Additionally, both the choices of separate-head attention over multi-head and slot-wise operations over the alternatives result in improved generalization, reduced model size, and fewer floating-point operations per second (flops). While the former supports the effectiveness of the separate-head attention bottleneck, the latter validates the benefits of the prior for shared separately-attendable concepts between the modalities enabled by Sparo.
@@ 5.4 Visualizations
To ease our interpretation of attended concepts, we choose concepts to visualize that have sharp attention in the text modality, attend to non-overlapping text tokens, and have cosine similarity more than 0.75 with the corresponding image slots. We present examples of these filtered concepts in Fig. 3. In this example, we see that Sparo is capable of attending to different concepts such as subject of the scene, the activity represented, and the surrounding location, for both vision and text. We provide additional visualizations in Sec. 0.B.7.
@@ 6 Discussion
Humans utilize a prior for compositionality to coherently represent different subsets of salient aspects of the world. We can enrich our comprehension by selectively attending to new aspects or specialize to a task by filtering out distracting ones. We introduce Sparo with a goal of imparting a similar prior to transformers in representation learning frameworks by partitioning encodings into separately-attended concepts. Although we see evidence of disentanglement through positive results of post-hoc concept selection and visualization of attention maps, we do not impose any explicit independence or disentanglement constraints. In our representation learning frameworks, the pressure for learning distinct concepts arises from the need to explain the factors of variation through the training objective with only the features that can be accumulated by a single head of attention with embedded queries and limited dimensionality. However, auxiliary training objectives to impose stronger distributional conditions on the learned concepts can be explored as promising future directions. Beyond pretraining, we highlight that more sophisticated post-hoc concept selection approaches than explored in our work, like using human interaction or set cover algorithms, can further improve the downstream utility of Sparo.
@@ Appendix 0.B Additional results
0.B.1 Hyperparameters $L$ and $V$
(a) ImageNet zero-shot accuracy.
(b) Number of model parameters.
Figure 0.B.1: Effect of varying $L$ and $V$ values for CLIP32+Sparo, and the embedding size for CLIP32 and CLIP32+GAP, when training on CC15M.
Sparo encodings are concatenations of $L$ separately-attended slots of $V$ dimensions each. We study the effect of $L$ and $V$ with CLIP32+Sparo trained on CC15M, comparing its ImageNet zero-shot accuracy and model size with that of CLIP32 and CLIP32+GAP in Fig. 0.B.1. Generally, performance is poor for very low values of $L$ and $V$, and improves when they are increased. Improvements remain monotonic with $L$ in our experimental range, but the returns diminish for larger values of $L$. For each increment of $L$, we add additional ${\bm{q}}_{l}$ and ${\bm{K}}_{l}$ parameters to the model, and considering the trade-off between model size and performance gains becomes vital. Unlike $L$, increasing $V$ beyond optimal values leads to a performance drop. With very large values of $V$, we produce low-rank slots that can potentially encode arbitrary harmful biases for downstream tasks.
0.B.2 Comparison with other bottlenecks
Table 0.B.1: Comparing zero-shot recognition, robustness, and SugarCrepe compositionality for CLIP models trained on Conceptual Captions with GAP, attentional pooling [41, 74], a linear (affine) information bottleneck, and a discrete bottleneck using Finite Discrete Tokens (FDT) [13].
Train Model ImageNet- Object Net Sugar Crepe
V1 V2 Sketch A R
CC3M CLIP32 (C) 0.109 0.091 0.053 0.025 0.154 0.064 0.578
$\mathcal{C}$+GAP 0.124 0.105 0.064 0.025 0.178 0.078 0.582
$\mathcal{C}$+AttnPool 0.114 0.092 0.045 0.026 0.149 0.070 0.589
$\mathcal{C}$+Sparo 0.132 0.113 0.072 0.027 0.190 0.083 0.596
CLIP16 (C) 0.141 0.122 0.068 0.033 0.177 0.080 0.611
$\mathcal{C}$+GAP 0.156 0.134 0.069 0.033 0.187 0.081 0.616
$\mathcal{C}$+AttnPool 0.161 0.138 0.074 0.036 0.199 0.090 0.599
$\mathcal{C}$+Sparo 0.170 0.140 0.088 0.035 0.221 0.098 0.625
CC12M CLIP32 (C) 0.304 0.257 0.201 0.060 0.405 0.160 0.691
$\mathcal{C}$+GAP 0.320 0.272 0.221 0.063 0.426 0.178 0.692
$\mathcal{C}$+AttnPool 0.290 0.240 0.170 0.046 0.357 0.165 0.679
$\mathcal{C}$+Sparo 0.344 0.293 0.251 0.069 0.468 0.200 0.693
CLIP16 (C) 0.361 0.311 0.249 0.091 0.467 0.218 0.697
$\mathcal{C}$+GAP 0.382 0.330 0.262 0.101 0.501 0.241 0.695
$\mathcal{C}$+AttnPool 0.357 0.305 0.226 0.085 0.428 0.226 0.685
$\mathcal{C}$+Sparo 0.406 0.350 0.298 0.113 0.559 0.268 0.723
CC15M CLIP32 (C) 0.329 0.279 0.232 0.070 0.435 0.178 0.687
$\mathcal{C}$+GAP 0.345 0.294 0.242 0.077 0.460 0.192 0.678
$\mathcal{C}$+AttnPool 0.312 0.262 0.190 0.065 0.374 0.168 0.671
$\mathcal{C}$+Sparo 0.370 0.313 0.271 0.083 0.510 0.213 0.710
CLIP16 (C) 0.384 0.337 0.268 0.105 0.503 0.238 0.699
$\mathcal{C}$+GAP 0.399 0.343 0.287 0.114 0.531 0.252 0.701
$\mathcal{C}$+AttnPool 0.387 0.329 0.248 0.100 0.469 0.233 0.685
$\mathcal{C}$+Sparo 0.437 0.378 0.317 0.145 0.579 0.279 0.730
Setup.
Linear information bottleneck details.
Discrete bottleneck details.
Results.
We present our results in Tab. 0.B.1. First, we see that a simple linear bottleneck performs worse than standard CLIP in a majority of settings, indicating that the mechanism behind the representational bottleneck is important. As expected, the discrete bottleneck widely outperforms or stays comparable to standard CLIP. However, note that global average pooling (GAP) continues to shine as a strong baseline, outperforming the discrete bottleneck in most settings. Finally, we verify that Sparo outperforms or remains comparable to the best baselines across all settings.
0.B.3 CLIP with ResNet encoders
Table 0.B.2: Comparing zero-shot recognition, robustness, and SugarCrepe compositionality for CLIP models trained on CC15M with ResNet-50 encoders.
Model Params ImageNet- Object Net Sugar Crepe
V1 V2 Sketch A R
[…]
We compare CLIP models with residual network (ResNet) [29] encoders, trained on CC15M with and without Sparo, in Tab. 0.B.2. Our findings are consistent with those observed with ViT encoders, demonstrating that Sparo improves model performance across SugarCrepe, ImageNet, and robustness benchmarks without increasing model size.
0.B.4 Fine-grained SugarCrepe performance
Table 0.B.3: Fine-grained SugarCrepe zero-shot accuracies for CLIP models trained on Conceptual Captions and LAION-400M.